OpenStudy (anonymous):

A cargo plane flew to the maintenance facility and back. It took one hour less time to get there than it did to get back. The average speed on the trip there was 220mph. The average speed on the way back was 200mph. How many hours did the trip there take?

OpenStudy (radar):

Ten hours d=speed(s) * times (t) Let t=the time to get back. then t-1 = time to get there. d is the distance d=220(t-1) d=200t d=d 220(t-1)=200t 220t-220=200t 20t=220 t=11 t-1=10 hours the time to get there.